From GlueXWiki
Jump to: navigation, search

August 19, 2010 FDC meeting

Tentative Agenda

  1. Production
  2. Electronics (Fernando)
    • update on first article PCBs
    • other
  3. Engineering
    • Cathode foil: sole source procurement status (Roger) form, PR 294196 questions
    • Wire stringing update (Bill)
    • Other: granite tables, update from UVA, returning the tent in 126 to HallA
  4. Full-scale prototype tests: update (Beni)
  5. Other: talks at the collaboration meeting


Participants: Ron Bartek and Carrol Jones (for the first part only), Fernando, Bill, Mark, Casey, David, Eugene, Tim, Beni, Simon, Christopher Stanislav, and Lubomir.


- Ron Bartek explained that the bids for the clean room came very high and listed his suggestions to reduce the price: 1) class 10,000 NOT OCCUPIED 3) humidity limits 35-55% 4) no outlets inside (at least initially) 5)max number of people from 10 to 6. 1) would mean tighter protocol (2) for entering/working in the clean room; we agreed to this change. 3) Bill suggested 35-50% humidity limits if the change from 55 to 50% is not critical for the cost. Carrol Jones gave us more details about the humidity control also at other places at JLab. 4) means that we can install outlets later. As for #5, maybe it is not so critical. In addition we discussed if the complete elimination of the exhaust is acceptable, if we need the vestibule as it is (Carrol explained us it is important). Ron Bartek will inform all of us tomorrow about the company's suggestions how to reduce the price and what dominates the price. This information will help greatly to continue this discussions.

- David made a list of materials and equipment needed for the FDC production, totaling $120K. We have to find a way to include this into the FDC budget. We discussed the main items: epoxy, tables, additional transfer ring, microscope, ... Tim suggested starting with smaller quantities and then buying/manufacturing depending on the needs. Eugene wanted to have these numbers spread over the years. Communicate with David to include other stuff in the list, if necessary.

- Mark and Casey prepared a list of possible defects for the FDC production. Eugene asked about the cathode soldering technique and the probability for failure. Bill is getting tomorrow the magic soldering tape (from 3M). We will have: G10 frame, cathode foil, magic tape and then the flex that has to be heated to 400F. Bill discussed the tooling needed for this technique: most likely we have to make it.


- Fernando presented Christopher Stanislav who is testing the PCBs. No problems with the first article PCBs. The Gerber files were modified to fix the coating problem: some of the vias were not coated. We will have all the PCBs ready by mid September. Bill: when we can have the first article PCBs to be used in the first wire frames? Fernando: probably end of September; working now on the procurement for the PCB stuffing.

- Fernando and Beni discussed the test of the prea-amps in discriminator configuration; Fernando has two pre-amps together with the converter to ECL ready. Fernando is also working on the procurement of the thin HV cables, that will be connected to the pins of the HV connector of the thick cables.


- Cathode foils: Roger communicates with the procurement about the sole source PR. They need more information about the foil production: why it is unique and can be done only at Allflex. They need independent estimate of the price. Fernando: suggested to ask Advanced Circuits for such estimate; the production is unique because of the large sizes of the foils and thin copper.

- no news about the wire stringing design; Bill is busy with other things.

- the two granite tables will arrive tomorrow at Blue Crab; Casey will coordinate this. We have to bring the position measurement system from UVA better by the end of next week. It is 4x4x2.5 ft and the best is to use a JLab van; Casey will talk to Tom about going to Charlottesville to bring the system. Hall A wants to use their tent in 126 for some short work; no objections. We will return the tent when we move to Blue Crab.

Full-scale prototype test

- Beni ran out of gas (Ar) last week and the testing was resumed today. He gave the fADC250 for firmware upgrade. It doesn't work yet and now switching back to the system without flash ADC.


- We will keep the same scheme for the FDC presentation at the next Collaboration meeting: Lubomir "FDC update", Beni "Full-scale prototype tests". Simon will talk about the FDC tracking in the following off-line section.

- Bill asked if 90/10 % Ar/CO2 will be our working gas mixture. Beni, Eugene, Simon, and Lubomir explained what are the pros and cons of low/high CO2 percentage: drift time fluctuations vs magnetic field effects. We will be in a position to decide about this at the beginning of the experiment.